# Visual Instruction Pretraining for Domain-Specific Foundation Models

### v2

@@ 4 Experiments
This section offers a rigorous assessment of ViTP’s finetuning capabilities on various downstream tasks from both remote-sensing and medical imaging domains. Dataset statistics and task specifications are summarized in Table 1. This section presents our experimental setup, a thorough analysis of the results, and an ablation study. We demonstrate ViTP’s superior performance compared to existing state-of-the-art methods and analyze its efficiency and robustness. The best score is indicated in bold, while the second-best score is underlined.
@@ 4.1 Remote Sensing
Table 2: Composition of the pretraining dataset for remote sensing.
       Dataset Size Sample Rate Tasks
  Mini-InternVL [38] 1394k 0.03 Caption,VQA,OCR
RSVQA [94] 100k 0.1 VQA
FIT_RS [98] 100k 0.1 VQA
GeoChat [61] 64k 2 VG
VRSBench [72] 38k 5 VG
RSVG [123] 5.5k 10 VG
DIOR-RSVG [162] 27k 8 VG
ISPRS_SAR [56] 1.5k 1 CLS
SAR_Sentinel-1&2 [131] 16k 1 CLS
VHM [109] 223k 1 Caption,VQA,CLS
LevirCCcaptions [87] 50k 0.5 Caption
GAIA [160] 33k 1 Caption
Million-AID [95] 920k 0.05 Caption,CLS
For the ViTP pretraining phase, we leverage a diverse set of publicly available image-text paired instruction datasets, adhering to the principles outlined in Section 3.3. The pretraining datasets are listed in Table 2. In the dataset recipe, Million-AID [95], GAIA [160], LevirCCcaptions [87], VHM [109], RSVQA [94], and FIT_RS [98] are large-scale remote sensing visual instruction datasets, providing diverse remote sensing scenes and visual tasks. ISPRS_SAR [56] and SAR Sentinel-1&2 [131] datasets specifically address the unique challenges of Synthetic Aperture Radar (SAR) data. We construct question-answer pairs related to object identification, scene interpretation, and attribute recognition in SAR images using these existing SAR classification datasets. This ensures modality alignment and task relevance for SAR-specific downstream applications. GeoChat [61], DIOR-RSVG [162], RSVG [123], and VRSBench [72] datasets primarily focus on visual grounding tasks within the remote sensing domain, providing rich instructional guidance for spatial reasoning, object attribute recognition, and target localization. The mini-InternVL [38] dataset contributes to general vision-language understanding capabilities, ensuring that ViTP retains broad applicability and does not overfit to highly specialized remote sensing patterns. The comprehensive pretraining dataset recipe enables ViTP to learn robust, instruction-aware visual representations that are highly relevant to diverse remote sensing tasks.
@@ 4.1.1 Object detection
Table 3 presents the object detection results on RGB remote sensing datasets. ViTP consistently achieves state-of-the-art performance across DIOR, DIOR-R, and DOTA-v2.0. For DOTA-v2.0, which is particularly challenging due to a large quantity of small objects and dense scenes, ViTP achieves a new state-of-the-art of 60.23 mAP. This significantly surpasses previous top performers like BillionFM (58.69), highlighting ViTP’s superior ability to handle complex spatial relationships and arbitrary object orientations, likely due to its instruction-following pretraining that fosters fine-grained localization capabilities. To be noticed, SkySense [41] pretraining requires significantly more computational resources, over $17\times$ the GPU hours of ViTP, ViTP still outperforms SkySense. These results underscore the effectiveness of ViTP’s instruction-following pretraining in learning robust and task-relevant visual representations for optical remote sensing imagery.
@@
Table 5: Semantic segmentation performance (mIoU %) on optical datasets.
@@ 4.2 Medical Imaging
Table 8: Composition of the pretraining dataset for medical imagery.
       Dataset Size Sample Rate Tasks
  Mini-InternVL [38] 1394k 0.02 Caption,VQA,OCR
GMAI-MMBench [158] 5k 20 VQA
Open-i [31] 7k 5 Caption
Huatuo-OA [15] 647k 0.1 Caption
Huatuo-VQA [15] 647k 0.5 VQA
PMC-OA [86] 1647k 0.05 Caption
PMC-VQA [169] 227k 0.5 VQA
OmniMedical [52] 89k 1 VQA
Quilt-1M [54] 723k 0.1 Caption
Quilt-Instruct [114] 147 1 VQA
4.2.1 Semantic Segmentation
@@ 5 Analysis and Ablation Study
Table 10: Validation of the data recipe. Performance on the RSAR Benchmark confirms the necessity of our comprehensive data curation strategy.
@@ 5.2 Impact of Pretraining Steps
Figure 5: Effect of Pretraining Duration. RSAR mAP improves with more pretraining steps before saturating at approximately 8k steps.
The duration of pre-training is a significant factor influencing downstream task performance. As depicted in Figure 5, the model’s mAP on the RSAR dataset generally improves with an increasing number of pre-training steps. This trend indicates that more extensive instruction tuning enables the model to learn progressively better feature representations. However, the performance curve begins to plateau around 8k steps, with no further improvements observed thereafter. Consequently, we set the number of ViTP pre-training steps to 8k by default for our main model, balancing performance with computational cost. This pre-training phase takes approximately 23 hours to complete 8k steps training on $8\times$ A40 GPUs, demonstrating the efficiency of our proposed method.
@@ 5.4 Impact of Language Model Size
This study investigates the influence of the Large Language Model’s (LLM) capacity on ViTP’s overall performance. As shown in Figure 7, as the LLM size increases, the downstream performance on the RSAR dataset steadily degrades. This phenomenon is analogous to the findings in Masked Autoencoders [50] pretrain paradigm, which highlight the importance of a lightweight decoder. It indicates that for the challenging SAR modality, a more powerful LLM might compensate for suboptimal visual features extracted by the Vision Transformer (ViT). This could reduce the optimization pressure on the ViT, preventing it from learning the most effective modality-specific representations.
@@ 5.5 Training Efficiency
A key advantage of the ViTP paradigm is its exceptional pretraining efficiency, which lowers the barrier for developing specialized foundation models. As detailed in Table 11, ViTP pretraining completes in one day on 8 A40 GPUs. This is dramatically faster than competing paradigms. For instance, ViTP is 2.6x faster than the efficient MIM-based method (Scale-MAE) and over 17x faster than the high-performing but computationally intensive contrastive learning method, SkySense.
@@
5.6 Data Efficiency
We investigate ViTP’s performance in low-data regimes by finetuning it on subsets of the RSAR training dataset, ranging from 2% to 20%. To fully assess the modeling potential of each method, we extend the finetuning schedule to 36 epochs for this specific study. As shown in Table 12, ViTP consistently outperforms SatMAE[27] (an MIM method) and RemoteCLIP [88] (a contrastive method) across all data fractions. ViTP’s advantage becomes more pronounced as the amount of training data decreases. For instance, with only 2% of the training data, ViTP achieves 46.98 mAP, substantially outperforming SatMAE (37.90 mAP) and RemoteCLIP (34.78 mAP). Notably, ViTP finetuned on just 20% of the data (67.07 mAP) achieves performance comparable to or exceeding many state-of-the-art methods trained on the full dataset (as shown in Table 4). This superior data efficiency suggests that the rich semantic priors acquired during instruction tuning enable the model to generalize more effectively from a limited number of examples.
5.7 Model Robustness
@@
6 Limitation and Future Work
Despite the strong performance and efficiency of ViTP, we acknowledge several limitations that open promising avenues for future research. A primary limitation is that ViTP’s success is intrinsically linked to the quality and diversity of the instruction-following dataset. As our experiments suggest, a carefully curated data recipe is important for achieving optimal performance. The current process of collecting, filtering, and balancing datasets for specialized domains like remote sensing or medical imaging requires considerable manual effort and deep domain expertise. Future work could focus on developing more automated or semi-automated strategies for generating high-quality, domain-specific instruction data. Leveraging large language models to synthesize diverse question-answer or description-grounding pairs from unlabeled images could be a promising direction to enhance the scalability and accessibility of our pretraining paradigm. Moreover, our current work focuses on pretraining 2D image-based Vision Transformers. Extending the ViTP framework to other data modalities, such as video and 3D point clouds, represents another exciting frontier. Adapting the instruction-following objective to these domains could unlock new capabilities in temporal reasoning and 3D spatial understanding.
7 Conclusion
In this paper, we addressed the underexplored, top-down pathway from high-level reasoning to low-level perception in vision foundation models. We introduced Visual insTruction Pretraining (ViTP), a novel paradigm that leverages the high-level semantic reasoning of a Vision-Language Model to directly guide the feature learning of a ViT backbone. Complemented by our Visual Robustness Learning (VRL), ViTP compels the vision model to learn rich, robust, and task-relevant representations from downstream-specific instructional data. Our extensive experiments on 16 challenging remote sensing and medical imaging benchmarks validated the effectiveness of the proposed ViTP method. ViTP not only established new state-of-the-art performance across a diverse range of downstream tasks but also proved to be remarkably computationally efficient. We believe this work opens a promising new avenue for a deeper integration of high-level reasoning into the core of visual feature learning.

### v3

@@ 4 Experiments
This section offers a rigorous assessment of ViTP’s finetuning capabilities on various downstream tasks from both remote-sensing and medical imaging domains. We present our experimental setup, a thorough analysis of the results, and an ablation study. We demonstrate ViTP’s superior performance compared to existing state-of-the-art methods and analyze its efficiency and robustness. The best score is indicated in bold, while the second-best score is underlined.
Implementation details, dataset descriptions, additional experiments on general-domain datasets, and extended analyzes are provided in the Supplementary Material.
@@ 4.1.1 Object detection
Table 1 presents the object detection results on RGB remote sensing datasets. ViTP consistently achieves state-of-the-art performance across DIOR, DIOR-R, and DOTA-v2.0. For DOTA-v2.0, which is particularly challenging due to a large quantity of small objects and dense scenes, ViTP achieves a new state-of-the-art of 60.23 mAP. This significantly surpasses previous top performers like BillionFM (58.69), highlighting ViTP’s superior ability to handle complex spatial relationships and arbitrary object orientations. To be noticed, SkySense [49] pretraining requires significantly more computational resources, over $17\times$ the GPU hours of ViTP, ViTP still outperforms SkySense. These results underscore the effectiveness of ViTP’s instruction-following pretraining in learning robust and task-relevant visual representations for optical remote sensing imagery.
@@
Table 3: Semantic segmentation (mIoU %) on optical datasets.
@@ 5 Analysis and Ablation Study
Table 7: Ablation of the data recipe. Performance on RSAR confirms the necessity of our comprehensive data curation strategy.
@@ 5.2 Impact of Pretraining Steps
Figure 5: Effect of Pretraining Duration. RSAR mAP improves with more pretraining steps before saturating at $\sim$8k steps.
The duration of pre-training is a significant factor influencing downstream task performance. As depicted in Figure 5, the model’s mAP on the RSAR dataset generally improves with an increasing number of pre-training steps. This trend indicates that more extensive instruction tuning enables the model to learn progressively better feature representations. However, the performance curve begins to plateau around 8k steps. Consequently, we set the number of ViTP pre-training steps to 8k by default for our main model, balancing performance with computational cost. This pre-training phase takes approximately 23 hours to complete 8k steps training on $8\times$ A40 GPUs, demonstrating the efficiency of our proposed method.
@@ 5.4 Impact of Language Model Size
This study investigates the influence of the LLM’s capacity on ViTP’s overall performance. As shown in Figure 7, as the LLM size increases, the downstream performance on the RSAR dataset steadily degrades. This phenomenon is analogous to the findings in Masked Autoencoders [55] pretrain paradigm, which highlight the importance of a lightweight decoder. It indicates that for the challenging SAR modality, a more powerful LLM might compensate for suboptimal visual features extracted by the ViT. This could reduce the optimization pressure on the ViT, preventing it from learning the most effective modality-specific representations.
@@ 5.5 Training Efficiency
A key advantage of the ViTP paradigm is its pretraining efficiency, which lowers the barrier for developing specialized foundation models. As detailed in Table 8, ViTP pretraining completes in one day on 8 A40 GPUs. This is dramatically faster than competing paradigms and achieves better downstream performance. For instance, ViTP is 2.6x faster than the efficient MIM-based method (Scale-MAE) and over 17x faster than the high-performing but computationally intensive contrastive learning method, SkySense.
6 Conclusion
This paper addressed the underexplored, top-down pathway from high-level understanding to low-level perception in vision foundation models. We introduced Visual insTruction Pretraining (ViTP), a novel paradigm that leverages the high-level semantic understanding of a Vision-Language Model to directly guide the feature learning of a ViT backbone. Complemented by our Visual Robustness Learning (VRL), ViTP compels the vision model to learn rich, robust, and task-relevant representations from downstream-specific instructional data. Our extensive experiments on 16 challenging remote sensing and medical imaging benchmarks validated the effectiveness of the proposed ViTP method. ViTP not only establishes new SOTA performance across a diverse range of downstream tasks but also proves remarkably computationally efficient. We believe this work opens a promising new avenue for a deeper integration of high-level understanding into the core of visual feature learning.
Appendix A More Related Work
A.1 Remote sensing Perception Foundation Model
A.2 Medical Images Perception Foundation Model
A.3 Continual Pretraining
Appendix B Implementation Details
Training a VLM from scratch conventionally requires a multi-stage pipeline: vision–language contrastive pretraining, projector alignment, and large-scale instruction tuning. To bypass the prohibitive cost of these initial phases, we instead bootstrap from a publicly available, high-capacity VLM. Specifically, we initialise our weights from InternVL-2.5 [24], which uses a custmized ViT-Large backbone [25] and a Qwen2 [3] language model. All pretraining are performed on 8× NVIDIA A40 (48 GB) GPUs with a global batch size of 128, and all downstream task finetunings are conducted on $8\times$ NVIDIA RTX3090 (24 GB) GPUs. For ViTP pretraining, we use the AdamW optimizer with a learning rate of 2e-5 and a cosine decay schedule. For remote sensing and medical image foundation models, we apply a schedual of 8,000 training steps, however for the general domain foundation model, we pretrain for 16,000 steps. To accelerate both pre-training and downstream finetuning, Flash-Attention [34] is integrated into every Vision Transformer self-attention layer, yielding extra throughput gain without altering convergence behaviour.
Appendix C Experimental Datasets
C.1 Pretraining Datasets
For the ViTP pretraining phase, we leverage a diverse set of publicly available image-text paired instruction datasets, adhering to the principles outlined in Section “Pretraining Dataset Recipe”.
C.1.1 General Domain
The pretraining datasets are listed in Table 9. In the dataset recipe, ShareGPT4V [21], DVQA [69], ChartQA [118], AI2D [71], DocVQA [119], and GeoQA+ [11] are diverse visual question answering (VQA) datasets, providing rich instructional data for complex reasoning, chart and document understanding, and geospatial question answering. The SynthDoG-EN [72] dataset specifically addresses the unique challenges of Optical Character Recognition (OCR) tasks. ShareGPT4V [21] also provides visual grounding (VG) data, enhancing spatial localization capabilities. GAIA [180] and Million-AID [110] contribute foundational captioning (Caption) and classification (CLS) tasks, ensuring the model develops robust general-purpose vision-language understanding. This comprehensive pretraining dataset recipe enables the model to learn robust, instruction-aware visual representations that are highly relevant to diverse general domain tasks.
Table 9: Composition of the pretraining dataset for general domain.
       Dataset Size Sample Rate Tasks
  ShareGPT4V [21] 767k 1 VQA,VG
DVQA [69] 200k 1 VQA
ChartQA [118] 18k 1 VQA
AI2D [71] 12k 1 VQA
DocVQA [119] 10k 1 VQA
GeoQA+ [11] 72k 1 VQA
SynthDoG-EN [72] 30k 1 OCR
GAIA [180] 33K 0.2 Caption
Million-AID [110] 920k 0.01 Caption,CLS
C.1.2 Remote Sensing Domain
The pretraining datasets are listed in Table 10. In the dataset recipe, Million-AID [110], GAIA [180], LevirCCcaptions [101], VHM [128], RSVQA [109], and FIT_RS [113] are large-scale remote sensing visual instruction datasets, providing diverse remote sensing scenes and visual tasks. ISPRS_SAR [65] and SAR Sentinel-1&2 [151] datasets specifically address the unique challenges of Synthetic Aperture Radar (SAR) data. We construct question-answer pairs related to object identification, scene interpretation, and attribute recognition in SAR images using these existing SAR classification datasets. This ensures modality alignment and task relevance for SAR-specific downstream applications. GeoChat [74], DIOR-RSVG [183], RSVG [143], and VRSBench [85] datasets primarily focus on visual grounding tasks within the remote sensing domain, providing rich instructional guidance for spatial understanding, object attribute recognition, and target localization. The mini-InternVL [46] dataset contributes to general vision-language understanding capabilities, ensuring that ViTP retains broad applicability and does not overfit to highly specialized remote sensing patterns. The comprehensive pretraining dataset recipe enables ViTP to learn robust, instruction-aware visual representations that are highly relevant to diverse remote sensing tasks.
Table 10: Composition of the pretraining dataset for remote sensing domain.
       Dataset Size Sample Rate Tasks
  Mini-InternVL [46] 1394k 0.03 Caption,VQA,OCR
RSVQA [109] 100k 0.1 VQA
FIT_RS [113] 100k 0.1 VQA
GeoChat [74] 64k 2 VG
VRSBench [85] 38k 5 VG
RSVG [143] 5.5k 10 VG
DIOR-RSVG [183] 27k 8 VG
ISPRS_SAR [65] 1.5k 1 CLS
SAR_Sentinel-1&2 [151] 16k 1 CLS
VHM [128] 223k 1 Caption,VQA,CLS
LevirCCcaptions [101] 50k 0.5 Caption
GAIA [180] 33k 1 Caption
Million-AID [110] 920k 0.05 Caption,CLS
Table 11: Composition of the pretraining dataset for medical imagery domain.
       Dataset Size Sample Rate Tasks
  Mini-InternVL [46] 1394k 0.02 Caption,VQA,OCR
GMAI-MMBench [178] 5k 20 VQA
Open-i [37] 7k 5 Caption
Huatuo-OA [18] 647k 0.1 Caption
Huatuo-VQA [18] 647k 0.5 VQA
PMC-OA [100] 1647k 0.05 Caption
PMC-VQA [190] 227k 0.5 VQA
OmniMedical [60] 89k 1 VQA
Quilt-1M [63] 723k 0.1 Caption
Quilt-Instruct [134] 147k 1 VQA
C.1.3 Medical Imagery Domain
C.2 Domain-Specific Finetuning Datasets
Table 12: Overview of downstream task datasets.
Task Dataset Modal Box Format Images Classes
  General Domain
  Detection and Segmentation COCO [99] RGB Polygons 118,287 80
COCO [99] RGB Hori. Box 118,287 80
ADE20K [198] RGB Mask 20,210 150
  Remote Sensing
  Object Detection DIOR [77] RGB Hori. Box 23,463 20
DIOR-R [30] RGB Ori. Box 23,463 20
DOTA-v2.0 [166] RGB Ori. Box 11,268 18
SARDet-100K [92] SAR Hori. Box 116,598 6
SSDD [188] SAR Hori. Box 1,160 1
RSAR [191] SAR Ori. Box 95,842 6
Semantic and Instance Segmentation iSAID [164] RGB Mask 2,806 15
LoveDA [161] RGB Mask 5,987 7
UAVid [114] RGB Mask 5,510 8
SSDD [188] SAR Polygons 1,160 1
Change Detection SVCD [75] RGB Mask 16,000 2
WHU [66] RGB Mask 11,456 2
LEVIR-CD [16] RGB Mask 10,192 2
S2Looking [135] RGB Mask 5,000 2
  Medical Image
  Medical Segmentation AMOS2022 [67] CT Mask 19,310 14
BraTS2021 [36] MRI Mask 76,467 3
CovidQUEx [144] X-ray Mask 5,826 2
C.2.1 Natural Image Detection and Segmentation
Object detection and segmentation in natural scenes are fundamental tasks in computer vision, serving as benchmarks for general-purpose visual understanding. Object detection involves localizing and classifying multiple objects within everyday photographs, while segmentation provides pixel-level masks for precise delineation. These tasks are crucial for applications ranging from autonomous driving to augmented reality. We evaluate ViTP’s versatility on standard natural image benchmarks using two complementary datasets: COCO [99] for object detection and instance segmentation, and ADE20K [198] for semantic segmentation. Following established protocols [55, 127], we adopt Mask R-CNN [57] as the default framework for COCO tasks with a 12-epoch training schedule and standard data augmentations. For ADE20K semantic segmentation, we employ UperNet [167] with a 160k iteration training strategy. We report standard COCO metrics (mAP@0.5:0.95) for detection and segmentation tasks on COCO, and mean Intersection over Union (mIoU) for ADE20K semantic segmentation.
C.2.2 Remote Sensing Object Detection
C.2.3 Remote Sensing Semantic Segmentation
C.2.4 Remote Sensing Change Detection
C.2.5 Medical Imaging Semantic Segmentation
Table 13: Comparison of pretraining paradigms for end-to-end finetuning. Results are shown for object detection ($AP_{box}$) and instance segmentation ($AP_{mask}$) on COCO val using Mask R-CNN (1x schedule), and semantic segmentation (mIoU) on ADE20K using UperNet (160k iterations). All models use a ViT-Large backbone.
Pretrain Paradigm Model COCO val ADE20K (mIoU)
(APbox) (APmask)
  Supervised IN. Cls. [40] 49.3 43.9 49.9
DeiT III  [152] 48.7 41.1 52.0
Contrastive Learning CLIP [129] 51.3 - 52.2
MoCov3 [26] 49.3 44.0 49.1
DINOv2 [127] 53.4 46.8 55.0
Masked Image Modelling BEiT [7] 53.3 47.1 53.3
MAE [55] 53.3 47.2 53.6
4M [121] 53.7 46.4 53.4
Understanding ViTP 53.9 46.7 55.8
Appendix D More Experiments
D.1 General Domains
To validate the efficacy of our proposed ViTP paradigm, we conducted comprehensive experiments against SOTA pretraining methods in two rigorous settings: end-to-end finetuning and frozen-backbone feature evaluation. All comparisons utilize a ViT-Large backbone for fair assessment.
As shown in Table 13, ViTP achieves good performance, posting top results on COCO object detection (APbox 53.9) and ADE20K semantic segmentation (mIoU 55.8). It significantly outperforms leading methods from all major pretraining paradigms, including supervised, contrastive learning, and Masked Image Modeling. On COCO instance segmentation, ViTP remains highly competitive (APmask 46.7). This consistent superiority highlights that the features learned via our top-down approach are highly effective for downstream adaptation.
To more directly assess the intrinsic quality of the learned visual representations, we performed semantic segmentation on ADE20K with the frozen Vision Transformer backbone. As detailed in Table 14, ViTP (L/16) achieves a 51.9 mIoU, surpassing other strong multimodal foundation models like AIMv2 [45] (51.5) and various SigLIP-based architectures.
ViTP’s leading performance in both end-to-end finetuning and frozen-feature evaluation settings provides compelling evidence for the efficacy of our top-down pretraining paradigm. This demonstrates that the VLM’s understanding-based objective implicitly trains the Vision Transformer to learn more powerful and generalizable representations, confirming that top-down understanding significantly benefits bottom-up perception.
Table 14: Semantic segmentation performance (mIoU) on ADE20K using an UperNet [167] architecture with a frozen ViT backbone.
   Model ViT Size ADE20K (mIoU)
  SigLIP [182] So/14 400M 35.2
SigLIP2 [69] So/16 400M 35.3
LLaVA-SigLIP [76] So/14 400M 39.9
COMP-SigLIP [22] So/14 400M 49.5
AIMv2 [45] L/14 300M 51.5
COMP-AIMv2 [22] L/14 300M 51.0
ViTP L/16 300M 51.9
@@ D.1 General Domains
D.2 Data Efficiency
We investigate ViTP’s performance in low-data regimes by finetuning it on subsets of the RSAR training dataset, ranging from 2% to 20%. To fully assess the modeling potential of each method, we extend the finetuning schedule to 36 epochs for this specific study. As shown in Table 15, ViTP consistently outperforms SatMAE[32] (an MIM method) and RemoteCLIP [102] (a contrastive method) across all data fractions. ViTP’s advantage becomes more pronounced as the amount of training data decreases. For instance, with only 2% of the training data, ViTP achieves 46.98 mAP, substantially outperforming SatMAE (37.90 mAP) and RemoteCLIP (34.78 mAP). Notably, ViTP finetuned on just 20% of the data (67.07 mAP) achieves performance comparable to or exceeding many state-of-the-art methods trained on the full dataset (as shown in main paper Table 2). This superior data efficiency suggests that the rich semantic priors acquired during instruction tuning enable the model to generalize more effectively from a limited number of examples.
D.3 Model Robustness
@@
Figure 8: Comparison of self-attention activation maps from the third ViT layer for a given query patch (red cross). The models were pretrained using ViTP (ours), CLIP, and DINOv2. Left: General nature images. Right: Abstract concept images that require a higher degree of human-like understanding.
Figure 9: Self-attention activation maps for a given query patch (red cross) from the ViTP pretrained Vision Transformer. Left: ViT is pretrained with remote sensing domain data. Right: ViT is pretrained with medical domain data.
Appendix E Visualization of Self-Attention Maps
To provide a qualitative analysis of the features learned by our ViTP backbone, we visualize its internal self-attention mechanisms. This comparison assesses how a ViT backbone, when pretrained with our Visual Instruction Pretraining (ViTP) , interprets semantic relationships within an image compared to backbones pretrained with leading visual foundation model pretraining methods: CLIP [129] and DINOv2 [127].
As shown in Figure 8 and Figure 9, we compute the self-attention activation maps for a specific query patch (marked with a red cross). These maps show which other image regions the model “attends to” when processing the information from that single query patch. In our analysis, we observed the ”artifacts phenomenon” (as discussed in previous works [35, 68]), where some layers produce diffuse attention. We found that the third layer of the Vision Transformer provides the clearest and most ideal representation of interpretable attention across all three methods, so we visualize the outputs exclusively from this layer.
In examples like the elephants, sheep, and giraffes, the query patch (red cross) prompts the ViTP backbone to attend to other, non-local parts of the image that are semantically related. This capability is even more pronounced in abstract images where the ”object” is a concept.
We also provide more visualization on domain specific images like remote sensing and medical images in Figure 9. This analysis confirms that our ViTP successfully imbue the ViT backbone with a high-level semantic understanding, enabling it to group visual information based on conceptual meaning rather than just low-level features.
Appendix F Limitation and Future Work
Despite the strong performance and efficiency of ViTP, we acknowledge several limitations that open promising avenues for future research. A primary limitation is that ViTP’s success is intrinsically linked to the quality and diversity of the instruction-following dataset. As our experiments suggest, a carefully curated data recipe is important for achieving optimal performance. The current process of collecting, filtering, and balancing datasets for specialized domains like remote sensing or medical imaging requires considerable manual effort and deep domain expertise. Future work could focus on developing more automated or semi-automated strategies for generating high-quality, domain-specific instruction data. Leveraging large language models to synthesize diverse question-answer or description-grounding pairs from unlabeled images could be a promising direction to enhance the scalability and accessibility of our pretraining paradigm. Moreover, our current work focuses on pretraining 2D image-based Vision Transformers. Extending the ViTP framework to other data modalities, such as video and 3D point clouds, represents another exciting frontier. Adapting the instruction-following objective to these domains could unlock new capabilities in temporal understanding and 3D spatial understanding.